# Simultaneous Uniformization and Algebraic Correspondences

### v1

@@ Abstract.
We prove a generalization of the Bers’ simultaneous uniformization theorem in the world of algebraic correspondences. More precisely, we construct algebraic correspondences that simultaneously uniformize a pair of non-homeomorphic genus zero orbifolds. We also present a complex-analytic realization of the Teichmüller space of a punctured sphere in the space of correspondences.
Both authors were supported by the Department of Atomic Energy, Government of India, under project no.12-R&D-TFR-5.01-0500 as also by an endowment of the Infosys Foundation. MM was also supported in part by a DST JC Bose Fellowship. SM was supported in part by SERB research project grant MTR/2022/000248.
@@ 1. Introduction
Algebraic correspondences (equivalently, finite-to-finite multi-valued maps with holomorphic local branches) on the Riemann sphere, viewed as dynamical systems, include rational dynamics and actions of Kleinian groups. This was observed by Fatou in the 1920s [Fat29]. Bullett and Penrose constructed the first examples of correspondences that combine the dynamics of quadratic rational maps and the modular group [BP94, BL20].
An orbit equivalence framework for combining/mating Fuchsian groups (or surfaces) with polynomials was developed in [MM23a]. The key idea of [MM23a] was to replace a Fuchsian group with a piecewise Möbius circle map (called a mateable map) that retains some of the key features of the group and is compatible with polynomial dynamics. Using the notion of mateable maps, combination theorems for these objects were proved. Principal examples of mateable maps associated with Fuchsian groups are given by Bowen-Series (Section 2.2) and higher Bowen-Series (Section 2.3) maps.
The above framework was extended to virtual orbit equivalences in [MM23c]. Specifically, one looks at a finite index subgroup $\Gamma_{0}^{\prime}$ of the original Fuchsian group $\Gamma_{0}$, and shows that the Bowen-Series map of $\Gamma_{0}^{\prime}$ (which is orbit equivalent to $\Gamma_{0}^{\prime}$) acting on the circle admits a quotient as a factor dynamical system that can be conformally mated with complex polynomials. Algebraic descriptions of these conformal matings were also given (initially under a real-symmetry assumption, which was later dropped in [LLM24]). This was used to construct correspondences on (possibly nodal) Riemann spheres. These correspondences are matings of complex polynomials and genus zero orbifolds.
@@
In this article, we show that this mating framework also has purely Teichmüller-theoretic consequences. Specifically, we establish the following generalization of the Bers’ Simultaneous Uniformization Theorem:
@@
The paper is organized as follows. Section 2 surveys the orbit equivalence and virtual orbit equivalence mating framework between genus zero orbifolds and polynomial dynamics leading up to the construction of algebraic correspondences arising as combinations of the corresponding Fuchsian groups and polynomials. In Section 3, we prove one of the main new results of this article. In particular, we show that the virtual orbit equivalence mating framework is ‘less demanding’ in that it allows one to manufacture a semi-global complex-analytic map of the Riemann sphere that combines a pair of topologically nonequivalent genus zero orbifolds. We also characterize this conformal mating as an explicit algebraic function, and then globalize it to obtain an algebraic correspondence that simultaneously uniformizes a pair of non-homeomorphic genus zero orbifolds. In the final Section 4, we use Theorem 1.1 to construct a holomorphic embedding of the Teichmüller space of a puncture sphere into the space of algebraic correspondences, each of which is generated by a Möbius involution and the local deck transformations of a rational map.
@@ 2.1. Mateable maps
    $A$ is orbit equivalent to $\Gamma$; i.e., the grand orbits of $A$ equal the orbits of $\Gamma$.
@@
    $A$ is an expansive covering map of degree $d$ greater than one.
@@
    $A$ is Markov; i.e., the maximal connected subsets of $\mathbb{S}^{1}$ on which $A$ is genuinely analytic give a Markov partition of $\mathbb{S}^{1}$ for $A$.
@@
Now, let $P$ be a a complex polynomial of degree $d>1$ (for our purposes, the qualitative features of $P$ will be similar to those of $z\mapsto z^{d}$).
@@ 2.2. Bowen-Series Maps
Let $\Sigma_{d}$ denote the $(d+1)-$punctured sphere. We construct a specific $2d-$sided ideal polygon in the unit disk symmetric about the $x-$axis. Let $11=z_{0},\cdots,z_{d}=-$ 1 denote the $2d$-th roots of unity on the upper semi-circle arranged counter-clockwise. The vertices of the ideal $2d$-gon are given by $z_{0},\cdots,z_{d},\overline{z_{1}},\cdots,\overline{z_{d-1}}$. The side-pairing transformations take the edge (bi-infinite geodesic) between $\overline{z_{i}},\overline{z_{i+1}}$ to the edge (bi-infinite geodesic) between $z_{i},z_{i+1}$ for all the middle edges; i.e., $i=1,\cdots,d-2$. The edge between $z_{0},\overline{z_{1}}$ is taken to the edge between $z_{0},z_{1}$. Similarly, the edge between $\overline{z_{d-1}}$ and $z_{d}$ is taken to the edge between $z_{d-1}$ and $z_{d}$. See [MM23b, Figure 2] for a diagram illustrating this situation. Let $\sigma_{1},\cdots,\sigma_{d}$ denote the associated Möbius transformations on the unit disk $\mathbb{D}$. The associated Bowen-Series map $A_{\mathrm{BS},d}:\mathbb{S}^{1}\to\mathbb{S}^{1}$ is the piecewise analytic map defined by
@@
[MM23a, Proposition 3.3, Theorem 3.7] Suppose that $d\geq 2$, so that $\Sigma_{d}$ has at least 3 punctures. Then for any marked group $\Gamma\in\mathrm{Teich}(\Sigma_{d})$, the Bowen-Series map $A_{\mathrm{BS},\Gamma}:\mathbb{S}^{1}\to\mathbb{S}^{1}$ is a degree $12d-$ 1 mateable map in the sense of Definition 2.1. In particular, $A_{\mathrm{BS},\Gamma}:\mathbb{S}^{1}\to\mathbb{S}^{1}$ is orbit equivalent to the Fuchsian group $\Gamma$.
@@ 2.3. Higher Bowen-Series Maps
Next, we fix a regular ideal 2d-gon $W$ as in Section 2.2. This will be the fundamental domain for a base Fuchsian group $\Gamma_{0}$ isomorphic to $\pi_{1}(\Sigma_{d})$, where (recall) $\Sigma_{d}$ denotes a $(d+1)-$punctured sphere. Concretely, assume that the ideal vertices of $W$ are the $2d$-th roots of unity. Let the vertices of $W$ on the lower semi-circle be numbered $1=1_{-}$, $2_{-}$cdots, $(d+1)_{-}=d+1$ in counterclockwise order. Let the vertices of $W$ on the upper semi-circle be numbered $11,2,\cdots,d+$ 1 in clockwise order (see Figure 2.1, see also [MM23a, Figure 3]).
@@
Figure 2.1. Fundamental domains for $A_{\Gamma_{0},\mathrm{aux}}$ and $A_{\Gamma_{0},\mathrm{hBS}}$: $4$ punctures.
@@
For two ideal boundary points $a,b$ of $R$, let ¿ $ab$ denote the maximal arc of $\mathbb{S}^{1}$ with endpoints at $a,b$ and no internal break-points; i.e., ¿ $ab$ has endpoints at the break-points $a,b$, and there do not exist any other ideal boundary points of $R$ in the arc. On ¿ $\ i_{-}(i+1)_{-}$ , define
@@
Note that $A_{\Gamma_{0},\mathrm{aux}}$ maps ¿ $\ i_{-}(i+1)_{-}$ to the closure of the complement of the arc ¿ $\ i(i+1)$
Next, for $i=1,\cdots,d$, and on each of the $d$ short arcs ¿ $\ \{i,j\}\{i,j+1\}$ for $i\leq j\leq i+d-1$ between $i,i+1$, define
@@
For $i\in\{2,\cdots,d\}$ and $11\leq j\leq i-$ 1, let $j=i-s$, so that $11\leq s\leq i-$ 1. Define
@@
on ¿ $\ \{i,j\}\{i,j+1\}$ Hence, $A_{\Gamma_{0},\mathrm{aux}}$ maps ¿ $\{i,j\}\{i,j+1\}$ to the long arc from $d-t$ to $d+1-t$ in a counterclockwise sense.
@@
Let $\widehat{A}_{\Gamma_{0},\mathrm{hBS}}$ be the canonical extension of $A_{\Gamma_{0},\mathrm{hBS}}$. It is easy to check that $\widehat{A}_{\Gamma_{0},\mathrm{hBS}}$ is a higher degree map without folding in the sense of Definition 2.6. The inner polygon for this higher degree map without folding is the ideal polygon with vertices at $11,2,\cdots,d+$ 1.
@@ 2.4. Virtually mateable maps
    $A$ is a factor of a possibly discontinuous circle endomorphism $\widetilde{A}$ such that the latter is orbit equivalent to a finite index subgroup of $\Gamma$.
@@
    $A$ is an expansive covering map of degree $d$ greater than one.
@@
    $A$ is virtually Markov; i.e., there exists $n\in{\mathbb{N}}$ such that the $n-$fold preimages of maximal connected subsets of $\mathbb{S}^{1}$ on which $A$ is genuinely analytic give a Markov partition of $\mathbb{S}^{1}$ for $A$.
@@ 2.5. Factor Bowen-Series maps
When $n\geq 3$, the Fuchsian group $\Gamma$ that uniformizes the surface $\Sigma$ admits a (closed) fundamental polygon $\Pi$, two of whose paired sides are given by the radial lines at angles 00 and $2\pi/n$. The remaining
@@
Moreover, when $n\geq 3$, the map $A_{\Sigma}^{\mathrm{fBS}}$ has $p$ critical points, each of multiplicity $n-1$. All these critical points are mapped to the same critical value. Finally, the factor Bowen-Series map $A_{\Sigma}^{\mathrm{fBS}}$ restricts to a self-homeomorphism of order two on $\partial\mathcal{H}$.
@@
[LLM24, Theorem 1.6] Let $P$ be a degree $d$ polynomial with connected Julia set. Suppose that $P$ is either
@@
[LLM24, Theorem 1.9][MM23c, Theorem B] Let $\Sigma\in\mathfrak{F}$ with the corresponding Fuchsian group $G$. Let $P$ be a degree $d(\Sigma)$ polynomial with connected Julia set which is either

### v2

@@ Abstract.
We prove a generalization of Bers’ simultaneous uniformization theorem in the world of algebraic correspondences. More precisely, we construct algebraic correspondences that simultaneously uniformize a pair of non-homeomorphic genus zero orbifolds. We also present a complex-analytic realization of the Teichmüller space of a punctured sphere in the space of correspondences.
Both authors were supported by the Department of Atomic Energy, Government of India, under project no.12-R&D-TFR-5.01-0500 as also by an endowment of the Infosys Foundation. $M$ was also supported in part by a DST JC Bose Fellowship. SM was supported in part by SERB research project grant MTR/2022/000248.
@@ 1. Introduction
Algebraic correspondences on the Riemann sphere $\widehat{\mathbb{C}}$, viewed as dynamical systems, include rational dynamics and actions of Kleinian groups. For the purposes of this paper, an algebraic correspondence on $\widehat{\mathbb{C}}$ is a finite-to-finite multi-valued map with holomorphic local branches. Equivalently, let $\pi_{1}$ and $\pi_{2}$ be the two canonical projections from $\widehat{\mathbb{C}}\times\widehat{\mathbb{C}}$ to the first and second factors respectively, and $Z$ be a codimension $1$ subvariety in $\widehat{\mathbb{C}}\times\widehat{\mathbb{C}}$ containing no fibers of $\pi_{1}$ and $\pi_{2}$; then $\pi_{2}\circ\pi_{1}^{-1}$ is said to be an algebraic correspondence from the first to the second factor. Algebraic correspondences, viewed as multi-valued maps, can be iterated in the obvious way (cf. [BP01]).
That rational dynamics and actions of Kleinian groups can be unified in the framework of correspondences was observed by Fatou in the 1920s [Fat29]. Bullett and Penrose constructed the first examples of correspondences that combine the dynamics of quadratic rational maps and the modular group [BP94, BL20].
Virtual orbit equivalence, and a combination framework
An orbit equivalence framework for combining/mating Fuchsian groups (or surfaces) with polynomials was developed in [MM23a]. The key idea of [MM23a] was to replace a Fuchsian group with a piecewise Möbius circle map (called a mateable map) that retains some of the key features of the group (namely, it is orbit equivalent to the group) and is compatible with polynomial dynamics (that is to say, it is topologically conjugate to $z^{d}$, for some $d\geq 2$). Using the notion of mateable maps, combination theorems for these objects were proved. Principal examples of mateable maps associated with Fuchsian groups are given by Bowen-Series (Section 2.2) and higher Bowen-Series (Section 2.3) maps of Fuchsian punctured sphere groups.
The above framework was extended to virtual orbit equivalences in [MM23c]. More precisely, one looks at a finite index subgroup $\Gamma_{0}^{\prime}$ of the original Fuchsian group $\Gamma_{0}$, and shows that the Bowen-Series map of $\Gamma_{0}^{\prime}$ (which is orbit equivalent to $\Gamma_{0}^{\prime}$) acting on the circle admits a factor dynamical system, called the factor Bowen-Series map. The restriction of a factor Bowen-Series map to the circle is topologically conjugate to $z^{d}$, for some $d\geq 2$, and hence it can be conformally mated with complex polynomials. Algebraic descriptions of these conformal matings were also given (initially under a real-symmetry assumption, which was later dropped in [LLM24]). This was used to construct correspondences on (possibly nodal) Riemann spheres. These correspondences are matings of complex polynomials and genus zero orbifolds.
@@
Statement of the main theorem
In this article, we show that this mating framework also has purely Teichmüller-theoretic consequences. Specifically, we establish the following generalization of the Bers’ Simultaneous Uniformization Theorem (cf. [Ber60]).
@@
A word about the general framework in which Theorem 1.1 belongs is in order. It is clear that a Kleinian group with a (symmetric) generating set consisting of $g_{1},\cdots,g_{k}$ is equivalent to the data of a reducible polynomial $P(x,y)=\Pi_{i}(y-g_{i}(x))$. The codimension $1$ variety $\{P(x,y)=0\}$ gives rise to an algebraic correspondence, whose branches are given by $g_{1},\cdots,g_{k}$.
However, the framework of correspondences is manifestly more flexible; it includes general rational functions and dynamical systems exhibiting features of rational maps and Fuchsian groups simultaneously (cf. [MM23c, LLM24]). In particular, uniqueness of a correspondence $\mathfrak{C}$ combining two orbifolds $\Sigma_{1},\Sigma_{2}$ in the sense of Theorem 1.1 is too much to hope for (see the higher Bowen-Series maps in [MM23a] for instance).
In the special case where $\Sigma_{1},\Sigma_{2}$ are homeomorphic hyperbolic orbifolds satisfying the hypotheses of Theorem 1.1, the theorem reduces to Bers’ Simultaneous Uniformization Theorem (by producing a generating set for the quasi-Fuchsian group that simultaneously uniformizes $\Sigma_{1},\Sigma_{2}$). It is in this sense that Theorem 1.1 is a generalization.
A sketch of the proof
Let us now briefly mention the main ideas that go into the proof of Theorem 1.1. The key source of flexibility in the proof is the virtual orbit equivalence mating framework mentioned above. Specifically, this mating framework does not require the surfaces $\Sigma_{1}$ and $\Sigma_{2}$, appearing in Theorem 1.1, to be homeomorphic; it is enough to assume that the factor Bowen-Series maps associated with $\Sigma_{1}$ and $\Sigma_{2}$ have equal degree. With this condition in place, we proceed to construct the conformal mating of the associated factor Bowen-Series maps (see Proposition 3.2).
We then inspect the conformal mating $F$ described in the previous paragraph on the boundary of its natural domain of definition, and observe that it is a boundary involution; i.e., an orientation-reversing involution on a piecewise real-analytic curve (see Proposition 3.4). Using an algebraic characterization of such meromorphic maps derived in [MM23c, LLM24], we write the conformal mating as an explicit algebraic function. Concretely, the conformal mating $F$ can be modeled by the standard involution $\eta(z)=1/z$ via a rational map $R$ of the Riemann sphere. The rest of the proof of Theorem 1.1 follows the strategy of [MM23c]. The algebraic description of $F$ allows us to promote it to an algebraic correspondence, which is defined as compositions of $\eta$ and the local deck transformations of $R$. Finally, arguments similar to the ones used in [MM23c] show that the correspondence uniformizes the surfaces $\Sigma_{1},\Sigma_{2}$ simultaneously, in a suitable sense (see Section 3.2 for details).
Organization of the paper
The paper is organized as follows. Section 2 surveys the orbit equivalence and virtual orbit equivalence mating framework between genus zero orbifolds and polynomial dynamics leading up to the construction of algebraic correspondences arising as combinations of the corresponding Fuchsian groups and polynomials. In Section 3, we prove one of the main new results of this article. In particular, we show that the virtual orbit equivalence mating framework is ‘less demanding’ in that it allows one to manufacture a semi-global complex-analytic map of the Riemann sphere that combines a pair of topologically nonequivalent genus zero orbifolds. Here we also characterize this conformal mating as an explicit algebraic function, and then globalize it to obtain the desired algebraic correspondences. In the final Section 4, we use Theorem 1.1 to construct a holomorphic embedding of the Teichmüller space of a puncture sphere into the space of algebraic correspondences, each of which is generated by a Möbius involution and the local deck transformations of a rational map (see Theorem 4.4).
@@ 2.1. Mateable maps
    AA is orbit equivalent to $\Gamma$; i.e., the grand orbits of $A$ equal the orbits of $\Gamma$.
@@
    AA is an expansive covering map of degree $d$ greater than one.
@@
    AA is Markov; i.e., the maximal connected subsets of $\mathbb{S}^{1}$ on which $A$ is genuinely analytic give a Markov partition of $\mathbb{S}^{1}$ for $A$.
@@
Now, let $P$ be a complex polynomial of degree $d>1$ (for our purposes, the qualitative features of $P$ will be similar to those of $z\mapsto z^{d}$).
@@ 2.2. Bowen-Series Maps
Let $\Sigma_{d}$ denote the $(d+1)$-punctured sphere. We construct a specific $2d$-sided ideal polygon in the unit disk symmetric about the $x$-axis. Let $1=z_{0},\cdots,z_{d}=-1$ denote the $2d$-th roots of unity on the upper semi-circle arranged counter-clockwise. The vertices of the ideal $2d$-gon are given by $z_{0},\cdots,z_{d},\overline{z_{1}},\cdots,\overline{z_{d-1}}$. The side-pairing transformations take the edge (bi-infinite geodesic) between $\overline{z_{i}},\overline{z_{i+1}}$ to the edge (bi-infinite geodesic) between $z_{i},z_{i+1}$ for all the middle edges; i.e., $i=1,\cdots,d-2$. The edge between $z_{0},\overline{z_{1}}$ is taken to the edge between $z_{0},z_{1}$. Similarly, the edge between $\overline{z_{d-1}}$ and $z_{d}$ is taken to the edge between $z_{d-1}$ and $z_{d}$. See [MM23b, Figure 2] for a diagram illustrating this situation. Let $\sigma_{1},\cdots,\sigma_{d}$ denote the associated Möbius transformations on the unit disk $\mathbb{D}$. The associated Bowen-Series map $A_{\mathrm{BS},d}:\mathbb{S}^{1}\to\mathbb{S}^{1}$ is the piecewise analytic map defined by
@@
[MM23a, Proposition 3.3, Theorem 3.7] Suppose that $d\geq 2$, so that $\Sigma_{d}$ has at least 3 punctures. Then for any marked group $\Gamma\in\mathrm{Teich}(\Sigma_{d})$, the Bowen-Series map $A_{\mathrm{BS},\Gamma}:\mathbb{S}^{1}\to\mathbb{S}^{1}$ is a degree 2d12d-1 mateable map in the sense of Definition 2.1. In particular, $A_{\mathrm{BS},\Gamma}:\mathbb{S}^{1}\to\mathbb{S}^{1}$ is orbit equivalent to the Fuchsian group $\Gamma$.
@@ 2.3. Higher Bowen-Series Maps
Next, we fix a regular ideal 2d-gon $W$ as in Section 2.2. This will be the fundamental domain for a base Fuchsian group $\Gamma_{0}$ isomorphic to $\pi_{1}(\Sigma_{d})$, where (recall) $\Sigma_{d}$ denotes a $(d+1)$-punctured sphere. Concretely, assume that the ideal vertices of $W$ are the $2d$-th roots of unity. Let the vertices of $W$ on the lower semi-circle be numbered $1=1_{-}$, $2_{-}$ $\cdots$, $(d+1)_{-}=d+1$ in counterclockwise order. Let the vertices of $W$ on the upper semi-circle be numbered $1,2,\cdots,d+1$ in clockwise order (see Figure 2.1, see also [MM23a, Figure 3]).
@@
Figure 2.1. Fundamental domains for $A_{\Gamma_{0},\mathrm{aux}}$ and $A_{\Gamma_{0},\mathrm{hBS}}$: 44 punctures.
@@
For two ideal boundary points $a,b$ of $R$, let ¿ $ab$ denote the maximal arc of $\mathbb{S}^{1}$ with endpoints at $a,b$ and no internal break-points; i.e., ¿ $ab$ has endpoints at the break-points $a,b$, and there do not exist any other ideal boundary points of $R$ in the arc. On ¿ $\ i_{-}(i+1)_{-}\$, define
@@
Note that $A_{\Gamma_{0},\mathrm{aux}}$ maps ¿ $\ i_{-}(i+1)_{-}\$to the closure of the complement of the arc ¿ $\ i(i+1)\$.
Next, for $i=1,\cdots,d$, and on each of the $d$ short arcs ¿ $\ \{i,j\}\{i,j+1\}\$for $i\leq j\leq i+d-1$ between $i,i+1$, define
@@
For $i\in\{2,\cdots,d\}$ and $1\leq j\leq i-1$, let j=isj=i-s, so that $1\leq s\leq i-1$. Define
@@
on ¿ $\ \{i,j\}\{i,j+1\}\$. Hence, $A_{\Gamma_{0},\mathrm{aux}}$ maps ¿ $\{i,j\}\{i,j+1\}$ to the long arc from dtd-t to d+1td+1-t in a counterclockwise sense.
@@
Let $\widehat{A}_{\Gamma_{0},\mathrm{hBS}}$ be the canonical extension of $A_{\Gamma_{0},\mathrm{hBS}}$. It is easy to check that $\widehat{A}_{\Gamma_{0},\mathrm{hBS}}$ is a higher degree map without folding in the sense of Definition 2.6. The inner polygon for this higher degree map without folding is the ideal polygon with vertices at $1,2,\cdots,d+1$.
@@ 2.4. Virtually mateable maps
    AA is a factor of a possibly discontinuous circle endomorphism $\widetilde{A}$ such that the latter is orbit equivalent to a finite index subgroup of $\Gamma$.
@@
    AA is an expansive covering map of degree $d$ greater than one.
@@
    AA is virtually Markov; i.e., there exists $n\in{\mathbb{N}}$ such that the $n$-fold preimages of maximal connected subsets of $\mathbb{S}^{1}$ on which $A$ is genuinely analytic give a Markov partition of $\mathbb{S}^{1}$ for $A$.
@@ 2.5. Factor Bowen-Series maps
When $n\geq 3$, the Fuchsian group $\Gamma$ that uniformizes the surface $\Sigma$ admits a (closed) fundamental polygon $\Pi$, two of whose paired sides are given by the radial lines at angles 0 and $2\pi/n$. The remaining
@@
Moreover, when $n\geq 3$, the map $A_{\Sigma}^{\mathrm{fBS}}$ has $p$ critical points, each of multiplicity n1n-1. All these critical points are mapped to the same critical value. Finally, the factor Bowen-Series map $A_{\Sigma}^{\mathrm{fBS}}$ restricts to a self-homeomorphism of order two on $\partial\mathcal{H}$.
@@
[LLM24, Theorem 1.6] Let $P$ be a degree $d$ polynomial with a connected Julia set. Suppose that $P$ is either
@@
[LLM24, Theorem 1.9][MM23c, Theorem B] Let $\Sigma\in\mathfrak{F}$ with the corresponding Fuchsian group $G$. Let $P$ be a degree $d(\Sigma)$ polynomial with a connected Julia set which is either